how do you get K in this equation 6(k+2)+(2k-5=22
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12569)

6(k+2)+(2k-5)=22
6k+12+2k-5=22
8k=22-12+5
8k=15
k=15/8 answer.
proof:
6(15/8+2)+2*15/8-5=22
90/8+12+30/8-5=22
120/8=22-12+5
120/8=15
15=15

Answer by stanbon(48575)

6(k+2)+(2k-5 = 22
6k + 12 + 2k -5 = 22
8k + 7 = 22
8k = 15
k = 15/8
==============
Cheers,
Stan H.